OpenStudy (anonymous):

Find the axis of symmetry of x=2. Also find the axis of symmetry of x^2+6x-1.

OpenStudy (anonymous):

f(x)=-2x^2+6x-1 graph( 300, 200, -6, 5, -10, 10, -2x^2 +6x -1) (graph 300x200 pixels, x from -6 to 5, y from -10 to 10, -2x^2 +6x -1). Vertex (1.5,3.5) Axis x=1.5 Points (0,-1) (1,3)

OpenStudy (anonymous):

hope that help

OpenStudy (campbell_st):

the general form of a quadratic is y = ax^2 + bx + c the axis of symmetry of the quadratic is found using x = -b/(2a) in your question a = 1 and b = 6 y=x^2 + 6x -1 the the axis is symmetry is x = -6/(2 x 1) x = -3

OpenStudy (campbell_st):

the axis of symmetry of x = 2 is y = 0... or the y-axis